Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-x – 8y = 43
4x + 32y = 172

Answer by ewatrrr(24785)


Hi
Solve using the elimination method
-x – 8y = 43 |multiplying 1st equation by 4 to eliminate x by adding equations
4x + 32y = 172
-4x – 32y = 172
4x + 32y = 172
0 ≠ 344 there is NO solution, these lines are parallel